Kevin has won a raffle and has the option of either receiving a single check for $125,000 now, or receiving a check for $50,000 each year for three years with the first check issued now. At what interest rate would James have to invest his winnings for him to be indifferent to the two options? Answer: 21.6%

I know the correct answer but I really don't understand how to get it. Please explain each step in your solution. Will up vote.
Answer #1

Present value of single check now=present value of annual payments for three years

125000=50000+50000/(1+r)^1+50000/(1+r)^2

50000/(1+r)^1+50000/(1+r)^2=75000

1/(1+r)^1+1/(1+r)^2=1.50

use trial and error method to find the value of r

after trying many values of r, we get r=21.6%

the above is answer..